OpenStudy (anonymous):

What is the solution to the equation 3 1 – 2x = 2 ?

OpenStudy (anonymous):

x ≈ –0.185 you have to take log of both sides

OpenStudy (anonymous):

click the equation button when you are putting in your problems makes it a lot easier to see

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

log(31 - 2x) = log(2x)??

OpenStudy (anonymous):

What is the solution to the equation 3^(1 + 2x) = 2?

OpenStudy (anonymous):

k so you can use the change of base and the inverse logarithm properties to solve this

OpenStudy (anonymous):

inverse property is \[\log_{a}a^x=x \]

OpenStudy (anonymous):

with this you can use 3 = a to eliminate the 3 and make life easier

OpenStudy (anonymous):

so \[\log_{3}3^(1-2x) \] becomes \[(1-2x) * \log_{3}3 \]

OpenStudy (anonymous):

that was the power rule

OpenStudy (anonymous):

\[\log_{3}3 = 1 \] because 3^1 = 3

OpenStudy (anonymous):

backtracking to your equation if you take the log on one side, you have to take the log on the other side as well so putting this all together, you have \[(1-2x)*1= \log_{3}2 \]

OpenStudy (anonymous):

for the \[\log_{3}2 \] part, you use the change of base formula, which is \[\log_{a}x=(\log_{b}x )/(\log_{b}a ) \]

OpenStudy (anonymous):

this gives you \[\log_{10}2 / \log_{10}3 \], which you can just plug into your calculator to get .6309297..

OpenStudy (anonymous):

so you have 1-2x=.6309297 (1-.6309297)/2 = .1845351232